using the knowledge of odd and even functions, find the Fourier series decomposition of

f(x)={2x,-π<x<π

{f(x+2π)


1
Expert's answer
2021-07-27T06:09:37-0400

For the odd function, the Fourier series is called the Fourier Sine series and is given by


"f_{odd}(x)=\displaystyle\sum_{n=1}^{\infin}b_n\sin nx"

where the Fourier coefficients are


"b_n=\dfrac{2}{\pi}\displaystyle\int_{0}^{\pi}f(x)\sin nxdx, n=1, 2, 3, ..."

"b_n=\dfrac{2}{\pi}\displaystyle\int_{0}^{\pi}2x\sin nxdx"

"\int 2x\sin nx dx"

"\int u dv=uv-\int v du"

"u=2x, du=2dx"

"dv=\sin nxdx, v=\int \sin nx dx=-\dfrac{1}{n}\cos nx"

"\int 2x\sin nx dx=-\dfrac{2x}{n}\cos nx+\int \dfrac{2}{n}\cos nxdx"

"=-\dfrac{2x}{n}\cos nx+\dfrac{2}{n^2}\sin nx+C"

"b_n=\dfrac{2}{\pi}\bigg[-\dfrac{2x}{n}\cos nx+\dfrac{2}{n^2}\sin nx\bigg]\begin{matrix}
 \pi \\
 0
\end{matrix}"

"=-\dfrac{4}{n}\cos \pi n=\dfrac{(-1)^{n+1}4}{n}"

"f(x)=\displaystyle\sum_{n=1}^{\infin}\dfrac{(-1)^{n+1}4}{n}\sin nx"